It's laundry time again, and you have a roll of $10 worth of quarters to take to the Wash-'n-Whirl. Each US quarter has width of 1.75 mm and a radius of 12 mm. What is the volume of the roll of quarters?

Answer by jim_thompson5910(35256)

There are 10/0.25 = 40 quarters total.

The total height of the stack of quarters is h = 40*1.75 = 70 mm

The volume is then...

V = pi*r^2*h

V = pi*12^2*70

V = 10080pi << ------- Exact Volume

V = 10080(3.14159)

V = 31667.2272 << ------- Approximate Volume


--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------